OpenStudy (anonymous):

1/3 + ½ = (1/x ) I know the answer is 6/5 but confused on how to get it. Help me out?

OpenStudy (mertsj):

1/3 + 1/2 = 2/6 + 3/6 =5/6 and so we have: \[\frac{5}{6}=\frac{1}{x}\]

OpenStudy (anonymous):

you are supposed to solve for x... @Mertsj

OpenStudy (mertsj):

Well then, do it. You now have two equal ratios. Cross multiply and solve

OpenStudy (anonymous):

ahh I see okay. Thanks! @Mertsj